OpenStudy (2yoututu11):

Solve the equation for y. 4x – 5y = 9

OpenStudy (anonymous):

y=9-4x/-5

OpenStudy (anonymous):

Subtract 4x from both sides -5y = 9-4x Divide bith sides by -5 y = (9-4x)/-5 y = (4x-9)/5

OpenStudy (anonymous):

Thank you thank you <33